The line in Rº given in Cartesian form by 5x+4y - 32=4, 5x+y-52=5you have to turn this into a vector form (x , y, z) = (?, ?,?)+t(?,?,?)
Answer #1

5+44-384 52 + y - 53=5 Eliminating & subtracting ♡ from O (54.5x) + 4y - y -33 +58 = 45 By tag = -1 Y = -1-27 3 Eliminating z옯 8 3 510 14 어이기 일 17 - = -3(3) (Dividing by 10 ) 3 경로) 니다 10 10x3 봇 IS - 이 꽃 t - lo ~15 X ★ + y스 105 킬 -157 (3) ( -12}lot 34